# FlashEdit: Decoupling Speed, Structure, and Semantics for Precise Image Editing

### v4

@@ Abstract
Figure 1: FlashEdit produces superior visual results for text-guided image editing, addressing background instability and semantic entanglement with an over $150\times$ speedup against DDIM [song2020denoising] + P2P [p2p].
footnotetext: * Equal contribution.footnotetext: † Corresponding author: Yulun Zhang, yulun100@gmail.com
@@ 1 Introduction
Text-guided image editing with diffusion models [instructpix2pix, dong2023prompt] has demonstrated remarkable capabilities, allowing users to perform complex semantic modifications with high fidelity. The standard methodology is built upon a two-stage inversion-denoising pipeline: an initial inversion process maps a source image to its corresponding noise latent, which is then progressively denoised to generate the edited output according to a target prompt [directInversion, masactrl]. The objective is to achieve high fidelity in both content preservation and target prompt alignment, which often necessitates a computationally intensive, multi-step process.
Recent research has pursued several distinct strategies to improve accuracy and speed. To tackle the latency of the multi-step denoising process, methods based on model distillation have been proposed to enable editing in a faster way [turboedit]. These approaches must carefully address challenges such as mismatched noise statistics and insufficient editing strength that arise when adapting multi-step frameworks to fast samplers [mokady2023null, miyake2025negative]. To improve edit precision and prevent semantic leakage into the background, another category of work modifies the model’s internal mechanisms, primarily by re-weighting or replacing attention maps to ensure the edit is spatially constrained [alphaedit, headrouter]. Recognizing that the final edit quality is highly dependent on the starting point, other approaches focus on refining the inversion technique itself [directInversion]. These methods aim to find a more accurate initial latent vector, with recent insights revealing that separating the objectives of content preservation and edit fidelity can yield significant performance gains and speedups [wang2025image].
@@
To address this multifaceted challenge, we introduce a novel editing methodology that establishes control at three progressively finer levels of granularity. At the foundational level, we tackle the macro-problem of temporal control. We propose a One-Step Inversion-and-Editing (OSIE) pipeline, built upon an ”Anchor-and-Refine” training strategy, which conquers the prohibitive latency of prior work and makes real-time interaction possible. With this temporal control established, we address the meso-level problem of spatial control. Our Background Shield (BG-Shield) mechanism provides structural integrity by performing a surgical intervention in the self-attention layers. It uses a background memory and foreground-core querying to create a hard separation between edited and unedited regions, guaranteeing background stability. Finally, with speed and structure secured, we target the micro-level problem of semantic control. We develop Sparsified Spatial Cross-Attention (SSCA), a refinement of the cross-attention mechanism that prunes irrelevant text tokens pre-softmax, ensuring the edit is guided by a clean, unambiguous semantic signal. Each component logically builds upon the last, forming a cohesive solution (Figure 1). Our main contributions can be summarized as follows:
@@
  • At the temporal level, we introduce the One-Step Inversion-and-Editing (OSIE) pipeline and its “Anchor-and-Refine” training strategy, which enables high-fidelity inversion for one-step diffusion models.
@@ 2.1 Diffusion Models
Recent advances in image synthesis have been largely driven by diffusion models [peebles2023scalable, kulikov2024flowedit], which have become a leading paradigm for generating high-fidelity images from text. The core mechanism involves an iterative denoising process that progressively refines a random noise vector into a coherent image conditioned on a text prompt. A landmark contribution in this area is Stable Diffusion  [stableDiffusion], a Latent Diffusion Model (LDM) [rombach2022high] that performs the computationally intensive denoising process in a lower-dimensional latent space, making the technology widely accessible. Parallel to this, alternative frameworks have emerged, such as Flow Matching models like Flux [flux2024]. Instead of an iterative refinement process, these models learn to map noise to an image via a more direct, straight-line trajectory, representing a different theoretical foundation for high-quality generative modeling.
To mitigate the high computational cost of these iterative models, various acceleration techniques have been proposed. Model quantization [li2024large, li2024fast, arbllm, pbs2p], cache mechanism [specee, specdiff], sparse attention [sparseAttn], pruning [specprunevla, yan2025recalkv], and distillation [hinton2015distilling, nguyen2025swiftedit] are general acceleration techniques for deep learning model. In diffusion models, specifically, one primary category is model quantization [dvdquant], which reduces memory footprint and computational load by converting full-precision model weights and activations into lower-bit representations. Another category involves cache mechanisms [liu2025reusing, deepcache, TaylorSeer2025], which enhance inference efficiency by exploiting temporal redundancy. These methods reuse intermediate features computed at earlier denoising steps to avoid redundant calculations in later steps. While effective in isolation, recent work like QuantCache [quantcache] demonstrates a unified framework can yield greater gains.
Figure 3: Illustration of our Background Shield (BG-Shield) mechanism. The top of the figure illustrates the problem of background inconsistency in standard editing, while the bottom details the pipeline of our method designed to solve it.
@@ 2.2 Editing Models
The task of editing real images with pre-trained generative models introduces the fundamental challenge of inversion: finding a latent representation that can faithfully reconstruct a given source image. This problem was first extensively studied in the context of Generative Adversarial Networks (GAN) Inversion [wang2021HFGI, zhu2020domain, zhu2016generative]. In comparison, DDIM Inversion [song2020denoising] provides a deterministic method to find a corresponding noise latent for a source image. Once this latent is obtained, various editing mechanisms are employed during the denoising process to apply the desired changes. A prominent family of methods focuses on attention control, where the cross-attention maps between text and image are manipulated. For example, to change a “photo of a red car” to a “blue car,” Prompt-to-Prompt [p2p] identifies the attention weights corresponding to the word “red” and replaces them with those for “blue,” preserving the attention for “car” and the background. Another powerful technique is feature injection, exemplified by Plug-and-Play (PnP) [pnp]. To preserve the identity of a subject, PnP injects the self-attention features—which encode structure and appearance—from the source image’s generation process into the edited one. A third approach is mask-based editing, where methods like DiffEdit [couairon2022diffedit] generate a mask indicating the region to be altered and then apply the denoising process only within that area. Despite these advances, a core challenge persists in perfectly disentangling the edited foreground from the unedited background.
@@ 3 Method
3.1 One-Step Inversion-and-Editing
Challenge: A Dual-Constraint Optimization Problem.  The task of learning an effective inversion mapping is fundamentally a dual-constraint optimization problem. The predicted noise latent, $\varepsilon_{inv}$, must simultaneously satisfy two competing objectives. The first is a fidelity constraint, requiring $\varepsilon_{inv}$ to encode sufficient information to perfectly reconstruct the source image. The second is a distributional constraint, requiring $\varepsilon_{inv}$ to adhere to the generator’s prior distribution, $\mathcal{N}(0,I)$, to ensure editability. While both constraints can be explicitly supervised when using synthetic data, the distributional constraint becomes non-trivial and unsupervised for real-world images where the ground-truth noise is unknown. Naively optimizing for fidelity alone causes a severe violation of the distributional constraint, leading to uneditable latents.
Motivation.  Our motivation is to design a training strategy that explicitly decouples and progressively solves these two constraints. We posit that the network must first learn to jointly satisfy both objectives in a fully-supervised setting before it can be adapted to handle the unsupervised nature of real-image inversion. This leads to our “Anchor-and-Refine” approach. The “Anchor” stage uses synthetic data to ground the network in a parameter space that respects both constraints. The ”Refine” stage then adapts this mapping to real images, where we introduce a powerful generative prior from a teacher model to act as a proxy for the now-unsupervised distributional constraint. This ensures that even for real images, fidelity is pursued without sacrificing editability.
Proposed Method.  Shown in Figure 2, our primary architectural modification is designed to resolve a fundamental tension in the inversion process. The inverted noise vector is typically burdened with two conflicting tasks: perfectly preserving the source image’s identity and remaining generic enough for subsequent editing. To decouple these roles, we introduce a dedicated visual adapter which provides the decoder $D$ with a direct visual information from the source image, inspired by [yin2024one, nguyen2025swiftedit, chen2025dove].
This way, the decoder’s output—the reconstructed latent $z^{\prime}$—becomes a function of three distinct inputs: the inverted noise $n$, the text condition $c_{t}$, and the explicit image features $c_{i}$. By directly supplying the visual identity via $c_{i}$, we liberate the noise vector $n$ from its strict reconstruction duty. It can now remain closer to a pure Gaussian distribution, drastically improving its malleability for downstream editing tasks.
Stage 1: Anchoring the Solution via Supervised Training.  The first stage aims to find a robust initialization, or ”anchor,” for the inversion network $I_{\theta}$. We use a synthetic dataset of $(\varepsilon_{gt},z_{0})$ tuples from the base generator $G$, which allows for direct and strong supervision. The training objective is twofold:
$\mathcal{L}_{\text{Stage1}}=\|z_{0}-\hat{z}_{0}\|_{2}^{2}+\|\varepsilon_{gt}-\varepsilon_{inv}\|_{2}^{2}.$ (1)
The $\|\varepsilon_{gt}-\varepsilon_{inv}\|_{2}^{2}$ term is critical in this stage. It constrains the network to a region of the loss landscape where its outputs naturally conform to the target distribution $\mathcal{N}(0,I)$. During this stage, we train both the inversion network $I_{\theta}$ and the image adapter. This teaches the adapter how to effectively provide visual priors that aid in reconstruction. This anchoring step prevents the network from converging to trivial solutions in the next stage.
@@ 3.1 One-Step Inversion-and-Editing
Stage 2: Refining with a Teacher-Student Objective.  With the network anchored, the second stage refines its mapping for the complexities of real-world images where the ground-truth noise $\varepsilon_{gt}$ is unknown. To prevent the distribution of $\varepsilon_{inv}$ from drifting, we introduce a regularization scheme framed as a teacher-student distillation process. We leverage a pre-trained ”teacher” model, $\phi$, to provide a dynamic, supervisory signal for our ”student” inversion network, $I_{\theta}$.
For each real image latent $z_{0}$, we first create a noisy version $z_{t}=\alpha_{t}z_{0}+\sigma_{t}\varepsilon_{inv}$ at a random timestep $t$. The teacher model $\phi$ then predicts the noise from this input, yielding a ”pseudo-ground-truth” target, $\varepsilon_{teacher}$:
$\varepsilon_{teacher}=\phi(\alpha_{t}z_{0}+\sigma_{t}\varepsilon_{inv},t,c).$ (2)
We then define a refinement loss, $\mathcal{L}_{\text{Refine}}$, that minimizes the L2 distance between our network’s output $\varepsilon_{inv}$ and the teacher’s prediction. Crucially, we treat the teacher’s output as a fixed target by applying a stop-gradient operator.
$\mathcal{L}_{\text{Refine}}=\mathbb{E}_{t}\left[w(t)\left\|\varepsilon_{inv}-\text{sg}\left(\varepsilon_{teacher}\right)\right\|_{2}^{2}\right],$ (3)
where $\text{sg}(\cdot)$ denotes the stop-gradient operation. This formulation turns the problem into a simple regression task where the student ($I_{\theta}$) is trained to produce a noise latent that the teacher ($\phi$) would have predicted. This distillation-style loss effectively regularizes the training, ensuring that for any given real image, the predicted noise $\varepsilon_{inv}$ is a solution that is not only perceptually accurate (as enforced by a parallel perceptual loss) but also highly plausible under the teacher’s learned world model.
In addition, we introduce a Cycle-Consistency Regularization loss, $\mathcal{L}_{\text{cycle}}$, to enforce that our inversion network $I_{\theta}$ and the one-step generator $G$ are functional inverses. We first compute the initial noise $\varepsilon_{inv}=I_{\theta}(z_{0},c_{src})$ and its reconstruction $\hat{z}_{0}=G(\varepsilon_{inv},c_{src})$. We then feed this reconstruction back into the inversion network to get a cycled noise, $\varepsilon_{cycle}=I_{\theta}(\hat{z}_{0},c_{src})$. The loss is defined as:
$\mathcal{L}_{\text{cycle}}=\left\|\varepsilon_{inv}-\varepsilon_{cycle}\right\|_{2}^{2}$ (4)
This loss ensures the noise from the real image is consistent with the noise from its own reconstruction.
@@ 3.2 Background Shield
DDIM [ddim] P2P [p2p] 17.87 208.80 219.88 71.14 25.01 22.44
NT-Inv [NT-Inv] P2P [p2p] 27.03 60.67 35.86 84.11 24.75 21.86
DDIM [ddim] MasaCtrl [masactrl] 22.17 106.62 86.97 79.67 23.96 21.16
Direct Inversion [directInversion] MasaCtrl [masactrl] 22.64 87.94 81.09 81.33 24.38 21.35
DDIM [ddim] P2P-Zero [pix2pixZero] 20.44 172.22 144.12 74.67 22.80 20.54
Direct Inversion [directInversion] P2P-Zero [pix2pixZero] 21.53 138.98 127.32 77.05 23.31 21.05
DDIM [ddim] PnP [pnp] 22.28 113.46 83.64 79.05 25.41 22.55
Direct Inversion [directInversion] PnP [pnp] 22.46 106.06 80.45 79.68 25.41 22.62
ReNoise(SDXL) [renoise] 20.85 176.84 51.78 72.44 24.41 21.88
TurboEdit [turboedit] 22.51 107.27 9.32 80.09 25.49 21.82
SwiftEdit [nguyen2025swiftedit] 23.31 71.54 6.18 82.25 25.56 21.91
[…]

### v5

@@ Abstract
Machine Learning, ICML
Figure 1: FlashEdit produces superior visual results for text-guided image editing, addressing background instability and semantic entanglement with an over $150\times$ speedup against DDIM (Song et al., 2020b) + P2P (Hertz et al., 2022).
@@ 1 Introduction
Text-guided image editing with diffusion models (Brooks et al., 2023; Dong et al., 2023) has demonstrated remarkable capabilities, allowing users to perform complex semantic modifications with high fidelity. The standard methodology is built upon a two-stage inversion-denoising pipeline: an initial inversion process maps a source image to its corresponding noise latent, which is then progressively denoised to generate the edited output according to a target prompt (Ju et al., 2023; Cao et al., 2023). The objective is to achieve high fidelity in both content preservation and target prompt alignment, which often necessitates a computationally intensive, multi-step process.
Recent research has pursued several distinct strategies to improve accuracy and speed. To tackle the latency of the multi-step denoising process, methods based on model distillation have been proposed to enable editing in a faster way (Deutch et al., 2024). These approaches must carefully address challenges such as mismatched noise statistics and insufficient editing strength that arise when adapting multi-step frameworks to fast samplers (Mokady et al., 2023a; Miyake et al., 2025). To improve edit precision and prevent semantic leakage into the background, another category of work modifies the model’s internal mechanisms, primarily by re-weighting or replacing attention maps to ensure the edit is spatially constrained (Fang et al., 2024; Xu et al., 2024). Recognizing that the final edit quality is highly dependent on the starting point, other approaches focus on refining the inversion technique itself (Ju et al., 2023). These methods aim to find a more accurate initial latent vector, with recent insights revealing that separating the objectives of content preservation and edit fidelity can yield significant performance gains and speedups (Wang et al., 2025b).
@@
To address this multifaceted challenge, we introduce a novel editing methodology that establishes control at three progressively finer levels of granularity. At the foundational level, we tackle the macro-problem of temporal control. At the foundational level, we tackle the macro-problem of manifold alignment. We propose a Cycle-Consistent One-Step Inversion (COSI) pipeline, built upon a bijective loop strategy, which conquers the prohibitive latency of prior work while maintaining high reconstruction fidelity. With this temporal control established, we address the meso-level problem of spatial control. Our Background Shield (BG-Shield) mechanism provides structural integrity by performing a surgical intervention in the self-attention layers. It uses a background memory and foreground-core querying to create a hard separation between edited and unedited regions, guaranteeing background stability. Finally, with speed and structure secured, we target the micro-level problem of semantic control. We develop Sparsified Spatial Cross-Attention (SSCA), a refinement of the cross-attention mechanism that prunes irrelevant text tokens pre-softmax, ensuring the edit is guided by a clean, unambiguous semantic signal. Each component logically builds upon the last, forming a cohesive solution (Figure 1). Our main contributions can be summarized as follows:
@@
  • At the temporal and geometric level, we introduce the Cycle-Consistent One-Step Inversion (COSI) pipeline, which establishes a bijective mapping between the image and noise latents to ensure alignment with the generator’s specific manifold.
@@ 2.1 Diffusion Models
Recent advances in image synthesis have been largely driven by diffusion models (Peebles and Xie, 2023; Kulikov et al., 2024), which have become a leading paradigm for generating high-fidelity images from text. The core mechanism involves an iterative denoising process that progressively refines a random noise vector into a coherent image conditioned on a text prompt. A landmark contribution in this area is Stable Diffusion  (Rombach et al., 2021), a Latent Diffusion Model (LDM) (Rombach et al., 2022) that performs the computationally intensive denoising process in a lower-dimensional latent space, making the technology widely accessible. Parallel to this, Flow Matching models like Flux (Labs, 2024) maps noise to an image via a more direct, straight-line trajectory, representing a different theoretical foundation for high-quality generative modeling.
To mitigate the high computational cost of these iterative models, various acceleration techniques have been proposed. Model quantization (Li et al., 2024a, b, 2025b; Yan et al., 2025b), cache mechanism (Xu et al., 2025; Pan et al., 2025), sparse attention (Li* et al., 2025), pruning (Wang et al., 2025a; Yan et al., 2025a), and distillation (Hinton et al., 2015; Nguyen et al., 2025) are general acceleration techniques for deep learning model. In diffusion models, specifically, one primary category is model quantization (Li et al., 2025a), which reduces memory footprint and computational load by converting full-precision model weights and activations into lower-bit representations. Post-Training Quantization (PTQ) is a particularly popular training-free approach in this domain. Another category involves cache mechanisms (Liu et al., 2025a; Ma et al., 2024; Liu et al., 2025b), which enhance inference efficiency by exploiting temporal redundancy. These methods reuse intermediate features computed at earlier denoising steps to avoid redundant calculations in later steps. While effective in isolation, recent work like QuantCache (Wu et al., 2025) demonstrates a unified framework can yield greater gains.
Figure 3: Overview of the Cycle-Consistent One-Step Inversion (COSI) framework. The framework consists of a Forward Path that maps the source image to a noise latent $z_{inv}$ for reconstruction, and a Consistency Path that re-inverts the reconstruction to enforce geometric symmetry. By minimizing the Cycle Loss between the initial noise $z_{inv}$ and the cycled noise $z_{cyc}$, COSI mathematically anchors the inversion to the generator’s manifold $\mathcal{M}_{G}$, preventing latent drift and ensuring stable, high-fidelity image editing.
@@ 2.2 Editing Models
The task of editing real images with pre-trained generative models introduces the fundamental challenge of inversion: finding a latent representation that can faithfully reconstruct a given source image. This problem was first extensively studied in the context of Generative Adversarial Networks (GAN) Inversion (Wang et al., 2022; Zhu et al., 2020, 2016). In comparison, DDIM Inversion (Song et al., 2020b) provides a deterministic method to find a corresponding noise latent for a source image. Once this latent is obtained, various editing mechanisms are employed during the denoising process to apply the desired changes. A prominent family of methods focuses on attention control, where the cross-attention maps between text and image are manipulated. For example, to change a “photo of a red car” to a “blue car,” Prompt-to-Prompt (Hertz et al., 2022) identifies the attention weights corresponding to the word “red” and replaces them with those for “blue,” preserving the attention for “car” and the background. Another powerful technique is feature injection, exemplified by Plug-and-Play (PnP) (Zhang et al., 2021). To preserve the identity of a subject, PnP injects the self-attention features—which encode structure and appearance—from the source image’s generation process into the edited one. This ensures that when changing a “photo of a person” to a “photo of a person smiling,” the facial identity remains consistent. A third approach is mask-based editing, where methods like DiffEdit (Couairon et al., 2022) generate a mask indicating the region to be altered and then apply the denoising process only within that area. Despite these advances, a core challenge persists in perfectly disentangling the edited foreground from the unedited background. Furthermore, most of these established techniques rely on computationally intensive multi-step inference, which severely hinders real-time interaction. While recent efforts have explored accelerating editing via model distillation (Garibi et al., 2024; Deutch et al., 2024; Nguyen et al., 2025), they often compromise the delicate balance between structural preservation and semantic alignment, failing to match the fidelity of iterative solvers.
@@ 3 Method
3.1 Cycle-Consistent One-Step Inversion
Challenge: Manifold Misalignment and Latent Drift.  Traditional inversion for multi-step diffusion relies on iterative refinement to push the noise latent into a region that perfectly reconstructs the source image. However, in one-step generative frameworks, the solution space is extremely narrow, and the generative trajectory is highly sensitive to the initial noise distribution. While distillation-based methods (Nguyen et al., 2025) attempt to map images to noise via a teacher’s guidance, they often suffer from manifold misalignment. In one-step models, the generator $G$ characterizes a precise, low-dimensional manifold $\mathcal{M}_{G}$ where the generative dynamics are significantly more compressed than multi-step counterparts. In such a compressed space, the model lacks the ”self-correction” capacity inherent in multi-step sampling. The predicted noise $\epsilon_{inv}$ from heuristic distillation may yield a visually plausible reconstruction $\hat{z}_{0}$ at the pixel level, but it frequently fails to reside on the exact manifold $\mathcal{M}_{G}$. This off-manifold noise introduces a hidden instability; when the latent is perturbed by a target prompt $c_{tgt}$ for editing, this initial ”manifold gap” is amplified by the one-step decoder. Without subsequent denoising steps to buffer this error, it manifests as latent drift. Consequently, the model exhibits catastrophic background flickering and structural collapses, as the starting point lacks the mathematical grounding required for stable, trajectory-consistent single-step denoising.
Motivation: From Mimicry to Bijective Manifold Anchoring.  Our key insight is that a robust inversion should not merely mimic a teacher’s noise prediction but must act as a functional, structural inverse of the generator $G$. In a one-step setting, where the diffusion ODE is approximated by a direct mapping, the relationship between the image latent $z_{0}$ and the noise $\epsilon$ should ideally approximate a bijective mapping within the generative subspace. Without such a constraint, the inversion mapping is ill-posed, often resulting in a ”many-to-one” collapse where distinct image details are erroneously mapped to identical noise regions, sacrificing both reconstructive fidelity and editability. We hypothesize that by enforcing a Cycle-Consistency Constraint, we can regularize the inversion network $I_{\theta}$ to find the ”canonical” noise latent. This canonical latent is mathematically anchored to the generator’s specific manifold, transforming the inversion-denoising path from an open-ended approximation into a self-contained, closed loop. By ensuring the inversion is ”aware” of the generator’s reconstruction logic, we preserve structural invariants even under aggressive semantic shifts. This anchoring effectively restricts the inverted noise to the high-probability region of the generator’s distribution, making the editing process inherently more predictable and robust.
Proposed Method: Cycle-Consistent Bijective Loop.  To resolve these issues, we propose COSI, which moves beyond simple distillation by integrating a self-supervised alignment objective. We define the inversion mapping as $\epsilon_{inv}=I_{\theta}(z_{0},c,c_{i})$, where $c$ is the text condition and $c_{i}$ is the visual adapter feature. While an initial foundational mapping is established using existing one-step training paradigms, our core contribution lies in bridging the manifold gap through the Inversion Cycle. Unlike previous works that treat the inversion network as a black-box regressor, COSI treats it as a symmetric partner to the generator, aiming for a mathematical ”closed loop.”
For any given reconstruction $\hat{z}_{0}=G(\epsilon_{inv},c)$, the cycle-consistency principle dictates that feeding this reconstructed image back into the inversion network must recover the original noise $\epsilon_{inv}$. We optimize $I_{\theta}$ by minimizing the deviation between the initial predicted noise and the noise re-extracted from the reconstruction:
$\mathcal{L}_{cyc}=\mathbb{E}_{z_{0},c}\left[\left\|\epsilon_{inv}-\text{sg}(I_{\theta}(G(\epsilon_{inv},c),c,c_{i}))\right\|_{2}^{2}\right],$ (1)
where $\text{sg}(\cdot)$ denotes the stop-gradient operation to ensure that the inversion network specifically learns to adapt to the generator’s fixed output characteristics. Unlike traditional distillation losses that rely on an external teacher $\phi$, $\mathcal{L}_{cyc}$ provides a geometry-aware signal that is intrinsically tied to the generator $G$. This objective forces $I_{\theta}$ to learn the precise inverse geometry of the one-step manifold by verifying its own predictions through the generator’s lens. By closing this loop, COSI ensures that the inverted noise resides within the high-fidelity region of the generative space, providing a robust, “invertible” starting point that maintains absolute background integrity even under complex semantic edits.
@@ 3.2 Background Shield
DDIM P2P 17.87 208.80 219.88 71.14 25.01 22.44
NT-Inv P2P 27.03 60.67 35.86 84.11 24.75 21.86
DDIM MasaCtrl 22.17 106.62 86.97 79.67 23.96 21.16
Direct Inversion MasaCtrl 22.64 87.94 81.09 81.33 24.38 21.35
DDIM P2P-Zero 20.44 172.22 144.12 74.67 22.80 20.54
Direct Inversion P2P-Zero 21.53 138.98 127.32 77.05 23.31 21.05
DDIM PnP 22.28 113.46 83.64 79.05 25.41 22.55
Direct Inversion PnP 22.46 106.06 80.45 79.68 25.41 22.62
ReNoise(SDXL) 20.85 176.84 51.78 72.44 24.41 21.88
TurboEdit 22.51 107.27 9.32 80.09 25.49 21.82
SwiftEdit 23.31 71.54 6.18 82.25 25.56 21.91
FlashEdit 25.33 62.25 4.28 83.32 25.51 22.29
FlashEdit(w/ GT masks) 25.28 62.43 4.33 83.02 25.58 22.32